Identify the solution(s) of the system of equations, if any.
x + 3y = 6
2x - 6y = 4

Answer by oscargut(2103)

x + 3y = 6
2x - 6y = 4
If we multiply by (2) the first eq
2x + 6y = 12
2x - 6y = 4
adding them
4x = 16
x = 4
y = 2/3


If you need more help with similar questions you can contact me now at:
mthman@gmail.com
Thanks